A string is 3.82 m long with a mass of 3.39 g. The string is held taut with a tension of 319 N applied to the string. A pulse is sent down the string. How long does it take the pulse to travel the 3.00 m of the string?


Round your answer to 5 decimal places


1
Expert's answer
2021-05-03T07:08:57-0400
"t=\frac{s}{v}=s\sqrt{\frac{m}{Fl}}\\t=3\sqrt{\frac{0.00339}{319\cdot3.82}}=0.005\ s"